OpenStudy (anonymous):

a=b(1/c-1/d) isolate c the problem says to divide both sides by b first

ganeshie8 (ganeshie8):

thats right, since we need to isolate c,

ganeshie8 (ganeshie8):

and we see that, b is multiplied there, so we perform the opposite operation - divide b,

ganeshie8 (ganeshie8):

\(a = b(\frac{1}{c} - \frac{1}{d})\) divide b both sides, \(\frac{a}{b} = \frac{1}{c} - \frac{1}{d}\)

OpenStudy (anonymous):

then I need to add 1/d to both sides

ganeshie8 (ganeshie8):

perfect !

OpenStudy (anonymous):

I just dont know what to do when I am left with 1/c

ganeshie8 (ganeshie8):

\(\frac{a}{b} + \frac{1}{d} = \frac{1}{c}\)

ganeshie8 (ganeshie8):

\(\frac{1}{\frac{a}{b} + \frac{1}{d}} \ = c\)

ganeshie8 (ganeshie8):

if you wanto simplify, u may simplify left hand side, by taking LCD...

ganeshie8 (ganeshie8):

but we are done wid what we asked, we isolated the c !!!!!!

OpenStudy (anonymous):

but i think i need to simplify the left side.

ganeshie8 (ganeshie8):

good :) simplify it by multiplying numerator & denominator wid the LCD = bd

OpenStudy (anonymous):

LCD, like the drugs? I dont know what it is.

ganeshie8 (ganeshie8):

\(\frac{bd*1}{bd*(\frac{a}{b} + \frac{1}{d})} \ = c \) \(\frac{bd}{ad+b} \ = c \)

ganeshie8 (ganeshie8):

lol, LCD = Least Common Denominator

OpenStudy (anonymous):

oh. that makes more sense. I was concerned that I was going to have to do drugs to pass my math classes

ganeshie8 (ganeshie8):

first thing pops up in mind for LCD is, Liquid Crystal Display hw come u get into drugs and all hmm

OpenStudy (anonymous):

Im clean! I promise!

ganeshie8 (ganeshie8):

lol you're funny ;p

ganeshie8 (ganeshie8):

good luck wid ur math exams :)

OpenStudy (anonymous):

It was on some online homework and I just got the answer back and it said I didnt have to simplify the fractions. so 1/(a/b+1/d) would have worked. If only I knew to begin with :/

OpenStudy (anonymous):

Thanks for your help

ganeshie8 (ganeshie8):

lol its okay yw !